Solve for x: 7[2-3(2-x)+2x]=3-2x
Answer by Alan3354(21612)

Solve for x: 7[2-3(2-x)+2x]=3-2x
-------------------
7[2-3(2-x)+2x]=3-2x
7[2 -6+3x + 2x] = 3-2x
7[-4 + 5x] = 3-2x
-28 + 35x = 3-2x
37x = 31
x = 31/37